The perimeter of a rectangle can be found by finding the total distance around the outside of the shape. The rectangle below has a length of 2x + 3 and a perimeter of 6x + 4. what is the answer to the problem?

Answer by ankor@dixie-net.com(22740)

The perimeter of a rectangle can be found by finding the total distance around the outside of the shape.
The rectangle below has a length of 2x + 3 and a perimeter of 6x + 4. what is the answer to the problem?
:
Assuming you wish to find the width (w) of the rectangle
:
2L + 2W = perimeter
we know L and the perimeter
2(2x + 3) + 2W = 6x + 4
4x + 6 + 2W = 6x + 4
2W = 6x - 4x + 4 - 6
2W = 2x - 2
divide both sides by 2
W = (x - 1); is the width of the rectangle